OpenStudy (ashes.boom):

Anyone online? I'm desperate for Algebra 2 help!!

OpenStudy (faiithhh):

same!! maybe we can help each other?

OpenStudy (ashes.boom):

Rational Equations in the Real World: Bake Sale The school is having a bake sale. The first cupcake you buy is $2, each additional cupcake is $1. Find the cost per cupcake for buying 10, 20, and 50. Write a rational equation for this scenario and then use it to solve. Show all work **IM NOT LOOKING FOR THE ANSWER I JUST NEED HELP UNDERSTANDING

OpenStudy (ashes.boom):

@Nnesha @IrishBoy123 @Frostbite

OpenStudy (faiithhh):

http://www.montereyinstitute.org/courses/Algebra1/COURSE_TEXT_RESOURCE/U11_L2_T1_text_final.html i found a website explaining rational equations if that helps at all?

OpenStudy (ashes.boom):

I understand rational expressions, i just dont understand how to get this problem represented as a rational expression. Ive been to that site earlier today and the time and rate problems dont help explain what i need. Thank you though

OpenStudy (faiithhh):

ohh okay. sorry i couldn't be of any help.

OpenStudy (ashes.boom):

@Jaynator495

OpenStudy (ashes.boom):

@zepdrix

zepdrix (zepdrix):

Hmm this is kind of weird. You been learning about sequences? Because that's how I'm thinking of the problem right now.

OpenStudy (ashes.boom):

not sequences. sorry. i have a function for it --> f(x) = 2 + (x-1) and i have tested it out and it works.... the problem is it has to be a rational expression.

OpenStudy (ashes.boom):

@puppyprinc3ss this is the question

zepdrix (zepdrix):

Well your expression certainly is rational :) Rational expressions look like this type of thing: \(\large\rm \dfrac{x+3}{x-2}\) That type of stuff. Your expression is: \(\large\rm \dfrac{2+(x-1)}{1}\) yay rational. Not sure what else you're trying to do with it :3

OpenStudy (ashes.boom):

so it counts as a rational equation even if there isnt an x in the denominator??

zepdrix (zepdrix):

We need to have some sort of polynomial in the numerator and denominator. The one in the denominator is a polynomial, it's just really simple, 1. Yaaa it should count :d

OpenStudy (ashes.boom):

do you think it matters if the graph doesnt look like all the other rational equations graphs

zepdrix (zepdrix):

Hmm what do all the other ones look like? D: Maybe I'm missing something..

OpenStudy (ashes.boom):

 and theres vertical and horizontal asymptotes that the graph doesnt cover

zepdrix (zepdrix):

Ya it's fine :) Just make sure you don't graph any of the negative numbers. We can't sell a negative amount of cupcakes.

OpenStudy (ashes.boom):

thank you so much for your help i really appreciate it

OpenStudy (sshayer):

you have to pay extra one dollar for first. so if you buy x cupcakes then cost of each caupcake\[=\frac{ x+1 }{ x }\]